OpenStudy (dumbsearch2):

Together, Tim and Alan can unload a truck 6 in minutes. When Tim works alone, it takes him 9 minutes longer than Alan to unload the truck. How long does it take each person to unload the truck alone?

OpenStudy (anonymous):

use the equation @hero provided earlier, in addition to the fact that tim takes 9 minutes longer than alan. this will give you a system of equations to work with

hero (hero):

Well, actually, you won't need to solve a system, but you will need to properly define variables.

OpenStudy (anonymous):

yeah if you define tim's time in terms of alan's time (or vice versa), that would suffice

hero (hero):

So to help you out, I will do most of the dirty work for you: Let T = the time it takes Tim to unload a truck alone A = the time it takes Alan to unload a truck alone Then the time it takes for them to unload a truck together is: \[\frac{T \times A}{T + A} = 6\] They give us that it takes Alan 9 minutes longer to unload the truck alone compared to the time it takes Tim so we can further define Alan's time as A = T + 9. We replace that in our formula to get: \[\frac{T (T + 9)}{T + (T + 9)} = 6\] If you do the math correctly, you will get the correct answer.

OpenStudy (dumbsearch2):

So 6's the answer right?:)

hero (hero):

You have to find the values of T and A bro. If you read the question again, you will realize this.

OpenStudy (dumbsearch2):

sry. :) thanks for you help! :)

hero (hero):

Do you have any clue how to solve it?

OpenStudy (dumbsearch2):

I get two solutions: t=-6 and t=9

OpenStudy (anonymous):

Which of those solutions makes sense in the context of this problem ?

hero (hero):

@dumbsearch2 : That's one of the questions to answer. The other question is, "Can you show the work that you've done to get that result?".

OpenStudy (dumbsearch2):

I get -6 and 9? Which solutions should I put where? Look at the attached stuff:

hero (hero):

I'm still looking for you to post your steps that show how you got those numbers. Also, only one of those numbers are correct; the other you have to discard since it doesn't make sense mathematically. Furthermore, you must realize you have only solved for T and still need to find A.

OpenStudy (dumbsearch2):

nope.... i still don't understand.... :( http://www.wolframalpha.com/input/?i=%28T%28T%2B9%29%29%2F%28T%2B%28T%2B9%29%29%3D6

OpenStudy (anonymous):

you solved for Tim's time, T. now ask yourself which of the following makes sense: Tim unloads a truck in -6 minutes (he finishes 6 minutes before he starts) Tim unloads a truck in 9 minutes (he finishes 9 minutes after he starts)

OpenStudy (anonymous):

keep in mind that math is a tool that models reality, but we have to keep an eye on the domain to make sure it corresponds with reality

hero (hero):

The fact that you're using Wolframalpha to do the computation for you shows that there's a lot you don't know, yet need to understand.

OpenStudy (dumbsearch2):

:(

OpenStudy (saifoo.khan):

My approach to this problem will be different.

OpenStudy (dumbsearch2):

@saifoo.khan what would it be?

hero (hero):

@saifoo.khan, it probably would, but trust me, even if you were to show the person your way, I doubt he would understand.

hero (hero):

This same user had trouble computing this: \[\frac{8 \times 56}{8 + 56}\]

OpenStudy (saifoo.khan):

O_O

hero (hero):

And that's with a calculator

hero (hero):

No disrespect towards you @dumbsearch2, but there's no need in hiding what you don't know.

hero (hero):

I'm not trying to be harsh, I'm just stating the reality that exists for you.

OpenStudy (dumbsearch2):

So the formula you provided is a franction? what do i have to solve for? i'll try...

OpenStudy (saifoo.khan):

I solved a similar problem here: http://openstudy.com/users/saifoo.khan#/updates/50fac4ffe4b022b322702102

OpenStudy (dumbsearch2):

The difference is "How long would it take them if they work together?" in this case it's "How long does it take each person to unload the truck alone?" so what would I change in the forumula or something? Thank you so much. :( Sorry for my ignorance. Just ignore me if it gets out-of-hand. :(

OpenStudy (saifoo.khan):

You just have to go in reverse order of this problem to solve your's.

OpenStudy (dumbsearch2):

So I want to solve for T, which is time? How do I solve the fraction (what Hero provided, the folula). I tried doing the distribuate property on the top of the franction but no luck. :(

OpenStudy (saifoo.khan):

Yes. You have to solve for T. Cross multiply. Get rid of : T+(T+9)

hero (hero):

Actually, you should probably simplify T + (T + 9) before doing anything. Afterwards, you'll want to multiply both sides by the expression in the denominator.

OpenStudy (dumbsearch2):

don't see how i could

hero (hero):

In order to simplify T + (T + 9), get rid of the parentheses. Then simplify T + T + 9 This will only work for the expression in the denominator. Don't worry about the expression in the numerator just yet.

OpenStudy (dumbsearch2):

I attempted to solve with the formula you provided, and I got to t^2 = 3t+54. What do I do from there, or is it incorrect? How should I go about solving this?

hero (hero):

Here's the deal @dumbsearch2, when you post the work you have done, you need to post all of your steps, that way we can properly verify your steps because if you make a mistake, we can go over that particular step you are having trouble with. So when you post what you have done, post everything, not just the step you are currently stuck on. It is important that we see what you are stuck on in addition to how you got there to begin with.

OpenStudy (dumbsearch2):

So, I simplified the denominator, just as you said. So I had this:T(T+9)/2T+9=6. Then, I multiplied the denominator on both sides. T(T+9)=6(2T+9). I distributed.T(T+9)=12T+54 Then...Now here is the part I am not sure about...I distribute the T: T^2+9T=12T+54. I moved the 9T to the right side: T^2=12T-9T+54. So now I have this: T^2=3T+54 I am not sure what to do next. Thanks so much for you help!!!

hero (hero):

Yes, that is correct. I guess I must have underestimated you because you have the correct steps thus far. At this point, you should recognize that what you need to do is place all terms on the left side to create a quadratic equation set equal to zero. Then solve the quadratic for T. Any time you solve a quadratic, you get two values. In this case, we only need one value for T, the other will need to be discarded as it is not relevant.

OpenStudy (dumbsearch2):

So I found 9 for Tim after solving the quadratic. But I am sort of confused about Alan. If I understand correctly, T=9+A. But if this is the case, then A is 0. Is this even possible? Thanks again for your help. :)

hero (hero):

Together, Tim and Alan can unload a truck 6 in minutes: \[\frac{T \times A}{T + A} = 6\] When Tim works alone, it takes him 9 minutes longer than Alan to unload the truck: T = A + 9 How long does it take each person to unload the truck alone? So what we have to do here is substitute the value of T into the working together equation: \[\frac{(A + 9) \times A}{2A + 9} = 6 \\A^2 + 9A = 6(2A + 9) \\A^2 + 9A = 12A + 54 \\A^2 - 3A - 54 = 0 \\A^2 - 9A + 6A - 54 = 0 \\A(A - 9) + 6(A - 9) = 0 \\(A - 9)(A + 6) = 0\] A = 9 or A = -6 -6 makes no sense for this type of problem. A = 9 is the correct value here. Now since Tim works 9 minutes longer, we have: T = A + 9 T = 9 + 9 T = 18 Thus Tim works for 18 minutes alone.

OpenStudy (dumbsearch2):

Thank you so much! :))) For your time! :) It works!!!! :)

OpenStudy (dumbsearch2):

Your the best person EVER! :D

hero (hero):

Sorry for the confusion bro. Now you see why it is important to correctly define variables.

OpenStudy (dumbsearch2):

Nope, now it's clear! I can't believe you helped me so much! :) Your so nice..

hero (hero):

Well, actually, initially, I was trying to help, but I spent the remainder of the time making up for the mistake I made while trying to explain it. I never want to leave knowing that all I did was confuse a student even more. It would not sit right with me.

OpenStudy (dumbsearch2):

Well thanks ever so much: the answer worked! :) It's correct, and I understand how you did it. :)

hero (hero):

Well, that's great. This should work for most if not all your working together problems.